OpenStudy (anonymous):

Can someone help e with this? I am having trouble understanding this subject. Simplify the quantity x plus 6 over 12, all over the quantity x minus 8 over 10.

OpenStudy (anonymous):

Here are my answers 5 x over quantity x minus 8 5 x over 2 open parentheses x minus 4 close parentheses 5 open parentheses x plus 6 close parentheses over 6 open parentheses x minus 8 close parentheses open parentheses x plus 6 close parentheses open parentheses x minus 8 close parentheses over 30

OpenStudy (anonymous):

here is the equation in number form

OpenStudy (anonymous):

This is the same as \[\frac{ x+6 }{ 12 }*\frac{ 10 }{ x-8 }\]

OpenStudy (anonymous):

I could cross multiply to make the 12 a 6, and the 10 a 5, but could I cross multiply the other two?

OpenStudy (anonymous):

help? @swagmaster47 @kelliegirl33

OpenStudy (anonymous):

(x+3/6)*(5/x-4) ?

OpenStudy (anonymous):

You dont need to cross multiply the way I set it up.

OpenStudy (anonymous):

((x+6)/12 ) / ((x-8)/10) = (x+6)/12 * 10/(x-8) = 5(x+6) / 6(x-8)

OpenStudy (anonymous):

Ok, I think I've got it. thanks

OpenStudy (anonymous):

one last question

OpenStudy (anonymous):

you both were a big help, who should I give the medal to?

OpenStudy (anonymous):

snackshack

OpenStudy (anonymous):

I can't award myself medals

OpenStudy (anonymous):

lol....too bad...give it to swagmaster...

OpenStudy (anonymous):

ok, thanks for your help!!!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

I gave you a medal anyways :)

OpenStudy (anonymous):

I was the one asking the question....?

OpenStudy (anonymous):

thnx